# Information-entropy-driven generation of material-agnostic datasets for machine-learning interatomic potentials

### v1

@@ 2.5 The accuracy/robustness trade-off of Rev-EM
The extreme diversity of Rev-EM could be expected to lead to unacceptably high errors on low-energy configurations that are the most thermodynamically-relevant at ambient conditions unless extremely high complexity MLIAPs are used. In order to mitigate this difficulty, ACE-2 was trained using an energy reweighting scheme that assigned 75% of the total weight to configurations within 2 eV/atom from the lowest energy configuration in the dataset, and 25% to the rest (see Sec. 5.5 for the details of the procedure). This reweighting exposes a hyperparameter that can be adjusted to tune the accuracy/robustness tradeoff. To demonstrate this, two other ACE potentials, referred to as ACE-1 and ACE-3, were trained with a high/low energy separation set to 1 eV/atom and 3eV/atom respectively (instead of 2 eV/atom for ACE-2). ACE-1 should intuitively lead to more accurate low-energy properties at the expense of higher errors in extreme conditions. In addition, the upper limit on forces has been decreased from 100 eV/Å  for the ACE-2 model to 50 eV/Å  for the ACE-1 model, again benefiting near-equilibrium structures, while the upper limit on the forces remains unchanged at 100 eV/Å  for ACE-3.
Model All Hull+1eV Hull+2eV
@@
Table 3: RMS test errors of the different W ACE models for energies and forces, respectively. The columns represent different energy ranges over which averages are taken: All over all configurations with energies within 20 eV/atom of the minimum energy structure in Rev-EM, Hull+1 eV over all configurations within 1 eV/atom of the minimum energy structure in Rev-EM, and Hull+2 eV over all configurations within 2 eV/atom of the minimum energy structure in Rev-EM.
@@
Encouragingly, we note that the performance of ACE-2 fitted to Rev-EM and tested on datasets available from the literature is quantitatively comparable to that of the MLIAPs originally trained on these datasets (see Table 1), suggesting that the accuracy/robustness tradeoff inherent to training to ultra-diverse datasets can be made very favorable. For example, the equivariant message-passing neural network NequiP [41, 33] that was trained on the Owen-W dataset yields MAEs of 85.13 meV/atom and 69.02 meV/Å, respectively for energies and forces, compared to MAEs of 37.76 meV/atom and 227.53 meV/Å  for ACE-2. While the energy errors for ACE-2 are about twice as small, its force errors are about three times as large, which is likely attributable to the relatively low weight assigned to forces compared to energies (90% vs 10%) in the ACE-2 training. Therefore, the much higher diversity of Rev-EM compared to Owen-W (see Fig. 3) does not inherently lead a to significant degradation in performance when training to Rev-EM once the contributions of each structures to the loss function are reweighted.
@@ 2.6 Rev-EM is material agnostic
As no material specific information is used during the generation of Rev-EM, it should in principle be equally applicable to other unary materials. To demonstrate this, MLIAPs were trained to six elements with different thermodynamics, namely fcc (Al), bcc (W), hcp (Be, Re and Os), and graphite (C) ground states. Except in the case of C (see details below), these datasets were generated through a uniform volume rescaling of all configurations in Rev-EM according to the density in their known ground states. In all cases except for C, MLIAPs were obtained using the ACE-2 settings described in Sec. 5.5.
Tab. 4 reports the test RMSE errors observed for these systems. Errors are significantly lower for Be and Al compared to Re and Os, which is consistent with previous observations that the behavior of late transition metals is especially difficult to capture [33], and with the expectation that absolute errors should scale roughly with the cohesive energy. Once again, the ACE-2 potential shows good performance when tested on the Owen dataset, which is available for Re and Os, exhibiting significantly lower errors than for testing on holdout from Rev-EM, as expected from the limited diversity of this targeted dataset. Similarly, the ACE-2 model tested on the Smith-Al dataset shows comparable performance to the ANI model trained in Ref. [14], albeit with larger force errors due to the relative energy/force importance in the training  [^1] We observed indications that the DFT calculations on two configurations in the Smith-Al dataset did not converge. These were removed from the analysis.. Note that the comparisons with literature datasets were based on published DFT data; inconsistencies between functionals and convergence parameters could slightly increase the measured test errors. The results once again show that MLIAPs trained to Rev-EM perform very well both on domain-expert and active-learning curated datasets.
@@
Once again the energy and force reweighting scheme ensures that low energy configurations are given more importance, and hence incur reasonable errors. Figs. 9 demonstrate that the ACE-2 potentials perform extremely well at predicting the properties of a broad range of different crystal structures, both in terms of absolute value and of relative ordering.
@@
(a) Equilibrium formation energy
(b) Equilibrium volume
(c) Bulk modulus
Figure 10: Performance of the Rev-EM ACE model (blue) on prototypical crystal structures for C, where A9 is the ground state graphite structure.
@@
As shown in Tab. 4, the model for C shows higher RMSEs on the test dataset compared to the rest of the elements, reflecting the unique and complex chemistry of C. Nonetheless, this model accurately predicts the energetics, equilibrium volumes, and bulk moduli of 23 different crystal structures, as shown in Fig. 10. We again emphasize that none of these structures were explicitly added to the training set, but that all structures were generated through the entropy maximization procedure. These encouraging results suggest that Rev-EM contains not only information regarding high energy structures, but also on thermodynamically-relevant low-energy configurations for a broad range of different materials.
@@ 3 Discussion
The unique and extreme diversity of Rev-EM makes it particularly challenging to capture using MLIAPs, especially when combined with a complex chemistry as in the case of C. As discussed above, a thorough exploration of the hyperparameter space of the MLIAPs (cutoff radius, energy reweighting, choice of the type of basis functions, etc.) is likely to significantly improve the accuracy of the Rev-EM models. In addition, simply increasing the complexity of the MLIAPs is also a promising avenue, although it incurs a corresponding computational cost. For example, Qamar et al. [42] demonstrated that increasing the complexity of ACE models results in a steady decrease of the observed errors for models of C. Repeating their experiment by increasing the number of basis functions from 387 to 488, causing an increase in adjustable parameters from 1094 to 1816, led to a decrease in test energy RMSE from 101.35 meV/atom to 84.12 meV/atom and of force RMSE from 750.67 meV/Å  to 645.76 meV/Å . As no error saturation was observed in Ref. [42] in this range of complexity, significant further reductions in errors can be expected through a continued increase in complexity of the potentials.
The initial indications provided here are therefore a promising step towards eventually the ultimate objective of uniform near-quantum accuracy over an extremely broad span of conditions. While the techniques to do so are being developed, hybrid approaches combining Rev-EM with an active-learning-driven refinement for specific materials appears to be a promising approach to tailoring the accuracy/transferability tradeoff to the needs of specific applications.
@@ 4 Conclusion
This study demonstrates that the curation of training datasets for MLIAPs can be recast as an optimization problem where the objective function is the information entropy of the features contained in the dataset. This reformulation of the problem enables the simple, systematic, and autonomous generation of ultra-diverse datasets. Extensive validation demonstrates that ACE potentials trained to Rev-EM are extremely robust up to large values of energies and forces, while capturing the properties of interest that are usually manually targeted during training set generation. Rev-EM is also shown to be transferable to a diverse set of other unary materials with different thermodynamics, including W, Al, Be, Re, Os, and C. While capturing the extended feature space sampled by Rev-EM is extremely challenging for MLIAPs, reweighting strategies, hyperparameter optimization, and increase in complexity of the MLIAPs offer multiple venues to achieve the best of both worlds where the tradeoff between accuracy and transferability is minimized.
@@ 5.5 ACE parameterization
The pacemaker [59, 10, 31] code was used for parameterization of nonlinear ACE [10] models for the Be, C, Al, W, Re and Os systems. An identical setup is used for all materials, without any explicit hyperparameter tuning. One of the material specific parameters is the cut-off radius which was set to 4.5 Å  for Be, 5.5 Å  for Al/C and 6.0 Å  for W/Re/Os. Another describes the parameterization of the short-range repulsion, given for all elements in the Supplementary Materials. The parameters of the short-range core repulsion (see Ref. [59]) have been estimated empirically without optimization.
@@
Author contributions AS refined the production code, generated the datasets, and trained the potentials. DP developed the research idea, secured funding, and wrote the initial version of the entropy maximization code. Both authors analyzed the results and wrote the manuscript.

### v2

@@ 2.5 The accuracy/robustness trade-off of Rev-EM
The extreme diversity of Rev-EM could be expected to lead to unacceptably high errors on low-energy configurations that are the most thermodynamically-relevant at ambient conditions unless extremely high complexity MLIAPs are used. In order to mitigate this difficulty, ACE-2 was trained using an energy reweighting scheme that assigned 75% of the total weight to configurations within 2 eV/atom from the lowest energy configuration in the dataset, and 25% to the rest (see Sec. 5.5 for the details of the procedure). This reweighting exposes a hyperparameter that can be adjusted to tune the accuracy/robustness tradeoff. To demonstrate this, two other ACE potentials, referred to as ACE-1 and ACE-3, were trained with a high/low energy separation set to 1 eV/atom and 3eV/atom respectively (instead of 2 eV/atom for ACE-2). ACE-1 should intuitively lead to more accurate low-energy properties at the expense of higher errors in extreme conditions. In addition, the upper limit on forces has been decreased from 100 eV/Å  for the ACE22-2- 2 model to 50 eV/Å  for the ACE11-1- 1 model, again benefiting near-equilibrium structures, while the upper limit on the forces remains unchanged at 100 eV/Å  for ACE-3.
Model All Emin+1eV Emin+2eV
@@
Table 3: RMS test errors of the different W ACE models for energies and forces, respectively. The columns represent different energy ranges over which averages are taken: All over all configurations with energies within 20 eV/atom of the minimum energy structure in Rev-EM, Emin+1 eV over all configurations within 1 eV/atom of the minimum energy (Emin) structure in Rev-EM, and Emin+2 eV over all configurations within 2 eV/atom of the minimum energy structure in Rev-EM.
@@
Encouragingly, we note that the performance of ACE-2 fitted to Rev-EM and tested on datasets available from the literature is quantitatively comparable to that of the MLIAPs originally trained on these datasets (see Table 1), suggesting that the accuracy/robustness tradeoff inherent to training to ultra-diverse datasets can be made very favorable, as was originally observed in Ref.[13]. For example, the equivariant message-passing neural network NequiP [49, 41] that was trained on the Owen-W dataset yields MAEs of 85.13 meV/atom and 69.02 meV/Å, respectively for energies and forces, compared to MAEs of 37.76 meV/atom and 227.53 meV/Å  for ACE-2. While the energy errors for ACE-2 are about twice as small, its force errors are about three times as large, which is likely attributable to the relatively low weight assigned to forces compared to energies (90% vs 10%) in the ACE-2 training. Therefore, the much higher diversity of Rev-EM compared to Owen-W (see Fig. 4) does not inherently lead a to significant degradation in performance when training to Rev-EM once the contributions of each structures to the loss function are reweighted.
@@ 2.6 Rev-EM is material agnostic
As no material specific information is used during the generation of Rev-EM, it should in principle be equally applicable to other unary materials. To demonstrate this, MLIAPs were trained to eight elements with different thermodynamics, namely fcc (Al), bcc (W), hcp (Be, Re and Os), graphite (C), trigonal A7 (Sb) and trigonal A8 (Te) ground states. Except in the case of C (see details below), these datasets were generated through a uniform volume rescaling of all configurations in Rev-EM according to the density in their known ground states. In all cases except for C, MLIAPs were obtained using the ACE-2 settings described in Sec. 5.5.
Tab. 4 reports the test RMSE errors observed for these systems. Errors are significantly lower for Be, Al, Sb and Te compared to Re and Os, which is consistent with previous observations that the behavior of late transition metals is especially difficult to capture [41], and with the expectation that absolute errors should scale roughly with the cohesive energy. Once again, the ACE-2 potential shows good performance when tested on the Owen dataset, which is available for Re and Os, exhibiting significantly lower errors than for testing on holdout from Rev-EM, as expected from the limited diversity of this targeted dataset. Similarly, the ACE-2 model tested on the Smith-Al dataset shows comparable performance to the ANI model trained in Ref. [18], albeit with larger force errors due to the relative energy/force importance in the training  [^1] We observed indications that the DFT calculations on two configurations in the Smith-Al dataset potentially did not converge. These were removed from our analysis.. Note that the comparisons with literature datasets were based on published DFT data; inconsistencies between functionals and convergence parameters could slightly increase the measured test errors. The results once again show that MLIAPs trained to Rev-EM perform very well both on domain-expert and active-learning curated datasets.
@@
Once again, the energy and force reweighting scheme ensures that low energy configurations are given more importance, and hence incur reasonable errors. Figs. 10 demonstrate that the ACE-2 potentials generally perform extremely well at predicting the properties of a broad range of different crystal structures, both in terms of absolute value and of relative ordering.
@@
(a) C (Equilibrium energies)
(b) C (Equilibrium volumes)
(c) C (Bulk moduli)
(d) Sb (Equilibrium energies)
(e) Sb (Equilibrium volumes)
(f) Sb (Bulk moduli)
(g) Te (Equilibrium energies)
(h) Te (Equilibrium volumes)
(i) Te (Bulk moduli)
Figure 11: Performance of the Rev-EM ACE model (blue) on prototypical crystal structures for C, Sb and Te.
@@
As shown in Tab. 4, the model for C shows higher RMSEs on the test dataset compared to the rest of the elements, reflecting the unique and complex chemistry of C. Nonetheless, this model accurately predicts the energetics, equilibrium volumes, and bulk moduli of 28 different crystal structures including BC8, $\beta$ Sn (A5), BCT5, hexagonal diamond (hex-dia) and fullerene (C60), as shown in Fig. 11. In addition to the prototypes, Table 5 reports the performance of the model on the (111), (100), and (110) diamond (A4) surfaces. ASE [50] was used to construct the (111) and the (100) surfaces, while the (110) reference surface structure is from Ref.[48]. The formation energies are computed by allowing only the relaxation of atomic positions until forces on all atoms are less 0.01 eV/atom. No surface reconstructions were taken into account. We again emphasize that none of these structures were explicitly added to the training set, but that all structures were generated through the entropy maximization procedure. These encouraging results suggest that Rev-EM contains not only information regarding high energy structures, but also on thermodynamically-relevant low-energy configurations for a broad range of different materials.
DFT ACE
(mJ/m2) (mJ/m2)
111 13.0 12.54
110 5.84 5.08
100 9.09 8.9
Table 5: Surface formation energies in diamond.
To demonstrate the dynamical stability of the C ACE model, MD simulations were performed in the NVT ensemble for 45 structures at temperatures ranging from 500 K to 10,500 K in steps of 1000 K. The structures were randomly selected from the Rev-EM dataset over densities in the range of 1.5-4.5 g/cc. Supercells of these structures were generated, resulting in cells containing anywhere between 400 to 900 atoms depending on the density. Each simulation was run successfully for a total of 150 ps with a time step of 0.25 fs, corresponding to a total of 67.5 ns of aggregate simulation time. All MD simulations were performed using the LAMMPS software package [51]. The configurations towards the end of the simulation were analyzed by computing their coordination numbers. At 3500 K, sp2 coordinated atoms make up more than 85% of the atoms for the low and intermediate density (1.5 to 3.0 g/cc) configurations. For configurations with densities in the range of 3.0-3.3 g/cc, a mixture of sp2 and sp3 hybridized atoms were found with the fraction of sp3 atoms being 50-70%. At densities close to 3.5 g/cc, sp3 hybridized atoms were found to constitute more than 80% of the atoms. As the volumes of the simulation cells are held fixed, configurations with densities above 4.0 g/cc result in pressures greater than 200 GPa. At these densities and pressures, the fraction of sp3 coordinated atoms is greater than 95% even at 4500 K. Snapshots extracted from three simulations generated at different densities are reported in the Supplementary Material (see Fig. S6). This demonstrates that the potential naturally produces very stable trajectories even in extreme conditions of temperature and pressure.
In addition, quench simulations were carried out as described in Refs. [14, 52]. 2000 atoms were randomly placed in a simple cubic cell at densities ranging from 1.5-3.5 g/cc. These supercells were melted at 10,000 K in the NVT ensemble and held for 10 ps, before quenching to 300 K by linearly decreasing the temperature at a rate of 100 K/ps, before holding the temperature constant for a further 13 ps, after which configurations were analyzed. The obtained fraction of sp3 bonds are in very good agreement with the DFT results [52], as shown in the Supplementary Material (see Fig. S7). Both the above simulations suggest that, the current C ACE model, accurately describes the governing crystal structures at all densities.
2.7 Comparison with existing datasets
MLIAP Dataset ANI [18] GAP [53] UNEP [40] Rev-EM
Al ANI [18] 2.57/60.28 233.89/143.95 209.64/55.29 73.34/77.99
GAP [53] 76.11/307.37 0.76/31.32 92.69/96.3 155.76/263.75
UNEP [40] 13.88/156.29 114.79/88.29 19.62/69.0 114.88/111.49
ACE-2 8.8/94.42 56.93/64.75 17.09/53.94 9.56/44.23
MLIAP Dataset DP [54] GAP [55] UNEP [40] Rev-EM
W DP [54] - 236.41/410.75 347.23/446.5 1280.09/2135.85
GAP [55] - 1.22/180.49 134.22/241.2 417.76/647.59
UNEP [40] - 61.74/370.52 27.77/246.57 520.85/795.87
ACE-2 - 70.87/252.13 44.37/263.54 60.95/433.6
Table 6: Performance of several MLIAPs for Al and W on existing datasets and on the Rev-EM dataset. The RMSE values are shown as energy/force in units (meV/atom)/(meV/Å ). Note that the entirety of the respective datasets were used to obtain these errors.
In order to compare the relative robustness of the potentials obtained from the the Rev-EM dataset, we tested our ACE models for Al, W, and C on multiple datasets available from literature. Configurations with energies above 10 eV/atom from the minimum energy structure and force magnitudes above 20 eV/Å have been removed from all the datasets for Al. Similarly, configurations with energies above 20 eV/atom from the minimum energy structure and force magnitudes above 100 eV/Å have been removed from all the datasets for W. The ANI [18], GAP [53] and Unified Neuroevolution Potential (UNEP) [40] models for Al were also cross-validated against all available datasets. For W, deep potential (DP) [54], GAP [55] and UNEP [40] models were again tested on the entire Rev-EM dataset and cross-validated on their respected training datasets when available. The results for Al and W are shown in Tab. 6. Note that the entire datasets were used to compute the errors, which might include both training and test data when assessing a model on its own dataset. The reported values for the ACE-2 models hence differ slightly from the values in Tabs. 1 and 4, where test errors are reported. For the same reason, the errors for MLIAPs tested on their own datasets may also differ from those reported in the original references.
The ACE-2 model for Al performs uniformly well on the ANI, GAP and UNEP datasets, as the errors are generally comparable to the training error on the Rev-EM dataset. On the other hand, the ANI, GAP, and UNEP models have very low test/train RMSEs when assessed on their own datasets, but cross-validation to other datasets shows much higher errors, indicating that different datasets likely contain mutually exclusive regions of configuration space. In all cases, the ACE-2 model trained to Rev-EM comes in second in terms of accuracy, only behind the matched models/dataset combinations. This indicates that the Rev-EM dataset produces models that are significantly more robust than alternative approaches. A broadly similar behavior is observed for W, where most potentials do significantly worse when tested on datasets other than the ones they were trained on, and especially when tested on the Rev-EM dataset, where almost 10-fold higher energy errors are observed when comparing other models to ACE-2. In contrast, the ACE-2 model trained to Rev-EM perform similarly when tested on different datasets, suggesting that Rev-EM is effectively a super-set of these.
MLIAP Dataset GAP-20 [56] GAP [25] SNAP [15] ANI-1xnr [16] ACE [14] Rev-EM
GAP-20U [17] 47.92/777.63 140.28/1113.17 - - 809.48/1232.56 1042.11/2846.62
GAP [25] 342.09/1200.09 95.93/589.93 - - 680.93/1440.19 687.36/2026.89
SNAP [15] 318.90/1659.66 631.15/1152.23 - - 2008.34/2115.11 1907.04/2322.9
ANI-1xnr [16] 360.34/794.76 243.35/679.56 - - 725.98/1037.11 1311.81/1523.53
ACE [14] 70.11/554.07 68.58/772.70 - - 86.67/655.43 675.04/5728.02
ACE (this work) 160.99/773.02 104.89/582.91 - - 214.20/924.11 105.27/948.75
Table 7: Performance of several MLIAPs for C on existing datasets and on the Rev-EM dataset. The RMSE values are shown as energy/force in units (meV/atom)/(meV/Å ). Note that the entirety of the respective datasets were used to obtain these errors.
Finally, Table 7 reports the performance of various MLIAPs for C: ACE [14], SNAP [15], GAP [56, 17, 25], and the ANI-1xnr [16], when tested on the Rev-EM dataset and cross-validated against the other datasets when available in the literature. In this case, configurations with energies above 10 eV/atom from the minimum energy structure and magnitude of forces above 40 eV/Å  have been removed from all the datasets. The energy and force RMSE for ACE (this work) is different from the values in Tab. 4 because the upper limit of force magnitude contained in the (test-only Rev-EM) dataset is increased to 40 eV/Å , while the training dataset had a maximum force magnitude of 25 eV/Å . Once again, irrespective of the underlying ML architecture and training sets, none of the MLIAPs show good transferability to the Rev-EM dataset, while the ACE model from this work shows comparable performance on all datasets.
As the cross-validation of MLIAPs on different datasets from literature is based on as-published DFT data, this includes inconsistencies due not only to different DFT settings (functionals/convergence parameters) but also due to different software packages used to generate their training datasets. To partially account for this this, we compute the RMSEs between $\Delta\mathrm{E}^{\mathrm{DFT}}$ and $\Delta\mathrm{E}^{\mathrm{MLIAP}}$ for all MLIAPs when cross-validated on other datasets. The $\Delta\mathrm{E}$ values are the energy differences from the respective minimum energy structure. However, absolute values are considered when comparing against matching MLIAPs/datasets. Forces are not affected by this correction.
@@ 3 Discussion
The results presented above demonstrate the the problem of developing truly robust MLIAPs that are transferable across a broad range of phases and structures remains an outstanding challenge even for "simple" unary systems. Indeed, many potentials previously developed by the community, sometimes over many years of iterative improvements, do show very high, "near quantum", accuracy for configurations similar to those in the training set but significantly higher errors (sometimes by a factor 10 to 100) on novel configurations. While problems with very well defined structural spaces benefit from the very high "local" accuracy provided by conventional approaches, many other problems where it is difficult to a priori delineate the relevant structural space, e.g., for systems under extreme conditions, can be expected to strongly benefit from approaches such as the ones presented here, even at the cost of a modest decrease in local accuracy.
Indeed, the extreme diversity of Rev-EM makes it particularly challenging to capture using MLIAPs, especially when combined with a complex chemistry as in the case of C. As discussed above, a thorough exploration of the hyperparameter space of the MLIAPs (cutoff radius, energy reweighting, choice of the type of basis functions, etc.) is likely to significantly improve the accuracy of the ACE models presented here. In addition, simply increasing the complexity of the MLIAPs is also a promising avenue, although it incurs a corresponding increase in computational cost. For example, Qamar et al. [14] demonstrated that increasing the complexity of ACE models results in a steady decrease of the observed errors for models of C. Repeating their experiment by increasing the number of basis functions from 387 to 488, causing an increase in adjustable parameters from 1094 to 1816, led to a decrease in test energy RMSE from 101.35 meV/atom to 84.12 meV/atom and of force RMSE from 750.67 meV/Å  to 645.76 meV/Å . As no error saturation was observed in Ref. [14] in this range of complexity, significant further reductions in errors can be expected through a continued increase in complexity of the potentials. It therefore appears possible to largely achieve both high accuracy and broad robustness, albeit at an increased computational cost.
@@ 4 Conclusion
This study demonstrates that the curation of training datasets for MLIAPs can be recast as an optimization problem where the objective function is the information entropy of the features contained in the dataset. This reformulation of the problem enables the simple, systematic, and autonomous generation of ultra-diverse datasets that distil a very high amount of information in a relatively small number of small configurations. Extensive validation demonstrates that ACE potentials trained to Rev-EM are extremely robust up to large values of energies and forces, while capturing the properties of interest that are usually manually targeted during training set generation. Rev-EM is also shown to be transferable to a diverse set of other unary materials with different thermodynamics, including Be, C, Al, Sb, Te, W, Re, and Os. While capturing the extended feature space sampled by Rev-EM is extremely challenging for MLIAPs, reweighting strategies, hyperparameter optimization, and increase in complexity of the MLIAPs offer multiple venues to achieve the best of both worlds where the tradeoff between accuracy and transferability is minimized. The initial indications provided here are therefore a promising step towards eventually meeting the ultimate objective of uniform near-quantum accuracy over an extremely broad span of conditions.
@@ 5.5 ACE parameterization
The pacemaker [71, 10, 39] code was used for parameterization of nonlinear ACE [10] models for the Be, C, Al, Sb, Te, W, Re and Os systems. An identical setup is used for all materials, without any explicit hyperparameter tuning. One of the material specific parameters is the cut-off radius which was set to 4.5 Å  for Be, 5.5 Å  for Al/C, 6.5 Å  for Sb, 7.5 Å  for Te and 6.0 Å  for W/Re/Os. Another describes the parameterization of the short-range repulsion, given for all elements in the Supplementary Materials. The parameters of the short-range core repulsion (see Ref. [71]) have been estimated empirically without optimization.
@@
Author contributions AS refined the production code, generated the datasets, trained the potentials, and analyzed the results. DP developed the research idea, secured funding, and wrote the initial version of the entropy maximization code. Both authors wrote the manuscript.